An investment earns 2.25% per annum compounded daily. How many years would be required for an investment to triple in value? Assume all years have 365 days.
I am struggling to grasp this one... please explain and show work!
Thanks so much for the help!

Answer by stanbon(75887)

An investment earns 2.25% per annum compounded daily. How many years would be required for an investment to triple in value? Assume all years have 365 days.
==========
A(t) = P(1+(r/n))^(nt)
-----
Your Problem
3P = P(1+(0.0225/365))(365t)
-----
(1+0.0000616)^(365t) = 3
---------
Take the log of both sides to get:
(365t) = log(3)/log(1.0000616)
----
365t = 17835
t = 48.86 years
=====================
Cheers,
Stan H.